### v1

@@ Abstract
Context: Observing the radio emission from exoplanets is among the most promising methods to detect their magnetic fields and a measurement of an exoplanetary magnetic field will help constrain the planet’s interior structure, star-planet interactions, atmospheric escape and dynamics, and habitability. Recently, circularly polarized bursty and slow emission from the Boötis ( Boo) exoplanetary system was tentatively detected using LOFAR (LOW-Frequency ARray) beamformed observations. If confirmed, this detection will be a major contribution to exoplanet science. However, follow-up observations are required to confirm this detection.
Aims: Here, we present such follow-up observations of the Boo system using LOFAR. These observations cover 70 of the orbital period of Boo b including the orbital phases of the previous tentative detections.
Methods: We used the BOREALIS pipeline to mitigate radio frequency interference and to search for bursty and slowing varying radio signals. BOREALIS was previously used to find the tentative radio signals from Boo.
Results: Our new observations do not show any signs of bursty or slow emission from the Boötis exoplanetary system.
Conclusions: The cause for our non-detection is currently degenerate. It is possible that the tentative radio signals were an unknown instrumental systematic or that we are observing variability in the planetary radio emission due to changes in its host star. More radio data (preferably multi-site) and ancillary observations (e.g. magnetic maps) are required to further investigate the potential radio emission from the Boötis exoplanetary system.
Follow-up LOFAR observations of the Boötis exoplanetary system
@@
A confirmed direct detection of exoplanetary radio emission has been elusive for decades. Observing the radio emission from exoplanets is among the most promising methods to detect their magnetic fields (Grießmeier 2015) and a measurement of an exoplanetary magnetic field will help constrain the planet’s interior structure, star-planet interactions, and atmospheric escape (e.g., Zarka et al. 2015; Grießmeier 2015; Lazio et al. 2016; Grießmeier 2017; Lazio 2018; Griessmeier 2018; Zarka 2018; Lazio et al. 2019). A magnetic field measurement can help also help break degeneracies in the mass-radius diagram (Lazio et al., 2019). Additionally, atmospheric dynamics may be influenced by the presence of a planetary magnetic field (e.g., Perna et al. 2010a; Rauscher & Menou 2013; Rogers & Komacek 2014; Hindle et al. 2021) and Ohmic dissipation might be one of the factors causing the anomalously large radii of hot Jupiters (e.g. Perna et al. 2010a, b; Knierim et al. 2022). Finally, a magnetic field might be one of the many properties needed on Earth-like exoplanets to sustain their habitability (e.g., Grießmeier et al. 2005, 2015, 2016; Lammer et al. 2009; Kasting 2010; Owen & Adams 2014; Lazio et al. 2010a, 2016; Meadows & Barnes 2018; McIntyre et al. 2019; Gronoff et al. 2020; Green et al. 2021b).
@@
Observations of auroral radio emissions produced via the Cyclotron Maser Instability (CMI) mechanism (Wu & Lee 1979; Zarka 1998; Treumann 2006) have been used to directly measure the magnetic field strengths of some of the magnetized planets in our Solar System (Burke & Franklin 1955; Zarka 1998). CMI radio signals are emitted at the local electron cyclotron frequency in the source region. Hence, the maximum gyrofrequency is directly proportional to the maximum magnetic field near the planetary surface. CMI radio emission is circularly polarized, beamed, and time-variable (e.g., Zarka 1998; Zarka et al. 2004).
@@
Many theoretical and observational studies have focused on studying exoplanetary radio emissions over the past few decades. The reviews by Zarka et al. (2015) and Grießmeier (2015, 2017) cover the subject in great detail. A large body of theoretical work111e.g., de Pater 2000; Farrell et al. 2004; Lazio et al. 2004; Stevens 2005; Grießmeier et al. 2005; Griessmeier2007PSS; Grießmeier et al. 2007b; Jardine & Collier Cameron 2008; Vidotto et al. 2010; Hess & Zarka 2011; Nichols 2011, 2012; Vidotto et al. 2012; Saur et al. 2013; See et al. 2015; Vidotto et al. 2015; Nichols & Milan 2016; Fujii et al. 2016; Vidotto & Donati 2017; Weber et al. 2017a, b, 2018; Lynch et al. 2018; Zarka 2018; Wang & Loeb 2019; Kavanagh et al. 2019; Vidotto et al. 2019; Shiohira et al. 2020; de Gasperin et al. 2020; Green et al. 2021b; Sciola et al. 2021; Narang et al. 2021; Hori 2021; Cendes et al. 2022 has been published since the foundational theoretical studies of Zarka et al. (1997), Farrell et al. (1999), Zarka et al. (2001), and Zarka (2007). Some of these studies aim to develop theoretical models that can predict the frequency and intensity of radio emissions observed from Earth. A variety of scaling laws have been developed to estimate the magnetic field strengths of exoplanets (e.g. Farrell et al. 1999; Sánchez-Lavega 2004; Grießmeier et al. 2007b; Christensen et al. 2009; Christensen 2010; Reiners & Christensen 2010). Likewise, the Radio-Magnetic Bode’s law is a common approach for estimating the radio emission intensity of these planets (e.g., Farrell et al. 1999; Lazio et al. 2004; Zarka 2007; Grießmeier et al. 2007b; Grießmeier 2017; Zarka et al. 2018; Zarka 2018).
@@
Over the past few decades, there have been many exoplanet radio observing campaigns, but none of these observations have resulted in an unambiguous detection222e.g., Yantis et al. 1977; Winglee et al. 1986; Zarka et al. 1997; Bastian et al. 2000; Farrell et al. 2003; Ryabov et al. 2004; Shiratori et al. 2006; George & Stevens 2007; Lazio & Farrell 2007; Smith et al. 2009; Lecavelier Des Etangs et al. 2009, 2011; Lazio et al. 2010a, b; Stroe et al. 2012; Hallinan et al. 2013; Murphy et al. 2015; Lynch et al. 2017; Turner et al. 2017a; O’Gorman et al. 2018; Lenc et al. 2018; Lynch et al. 2018; Route 2019; Narang et al. 2020; de Gasperin et al. 2020; Green & Madhusudhan 2021a; Turner et al. 2021; Narang et al. 2021; Cendes et al. 2022; Narang 2022; Narang et al. 2023a, b; Route & Wolszczan 2023; Turner et al. 2023; Shiohira et al. 2024; Narang et al. 2024. There have been a few tentative detections (Lecavelier des Etangs et al., 2013; Sirothia et al., 2014; Vasylieva, 2015; Bastian et al., 2018), but none of them have been confirmed by follow-up observations. The many degenerate reasons for the non-detections are discussed in great detail in the review articles Zarka et al. (2015) and Grießmeier (2015, 2017). Some reasons for the non-detections could be that the observations were not sensitive enough (e.g., Bastian et al. 2000), there was no emission at the observed frequencies due to a low magnetic field strength (e.g. Hallinan et al. 2013), Earth was outside the CMI beaming pattern (Hess & Zarka 2011; Ashtari et al. 2022), and/or CMI mechanism was not operating (e.g., plasma frequency of the planet’s ionosphere was greater than the cyclotron frequency; Grießmeier et al. 2007b; Weber et al. 2017a, b; Lamy et al. 2018; Weber et al. 2018). By contrast, advancement has been made in detecting free-floating planets near the brown dwarf boundary (Kao et al. 2016, 2018) and stellar emission potentially created by star-planet interactions (e.g., Vedantham et al. 2020; Callingham et al. 2021; Pérez-Torres et al. 2021; Pineda & Villadsen 2023; Trigilio et al. 2023; Blanco-Pozo et al. 2023).
@@ 2 Observations
All of the observations presented here were taken with the Low Band Antenna (LBA) of LOFAR (van Haarlem et al. 2013) using beamformed mode (Stappers et al. 2011). We obtained 8 exoplanet observations (each 8 hours long) for a total of 64 hours, plus 8 observations of the pulsar B0809+74 totaling 82 minutes. Similar to our previous studies (Turner et al. 2017a, T19, T21, Turner et al. 2023), the pulsar observations are used as a calibrator to study systematics in the data and to ensure the reliability of the processing. An example OFF-beam dynamic spectrum is shown in Figure 2. The dates and times of the observations can be found in Table 1. The setup of the observations can be found in Table 2; it is similar to that in T21. Most importantly, we ensure we cover the 15–30 MHz range where T21 detected their tentative signals. We only present the Stokes-V333Similar to T21, we only analyze as defined in equations 9–11 of Section 4.2.1 in T19. in this paper since the tentative detections were seen in Stokes-V (T21).
@@ 3 Data pipeline
We performed the post-processing of the LOFAR data exactly as in T21. A very detailed description of the burst and slow emission post-processing observables can be found in T19 and T21. A brief summary is given in the following paragraph. The post-processing is performed on the absolute value of the corrected Stokes-V data (as defined in T19). For burst emission, we only use the Q2 and Q4a-f observable quantities. The Q2 observable consists of one time series at a time resolution of 1 second per beam (ON, OFF1, OFF2, and OFF3). It is obtained by high-pass filtering the dynamic spectrum. To search for faint emission, the high-pass filtered dynamic spectrum is integrated over a large frequency range (e.g., 10 MHz) to produce a time series. Q2 can be represented by a “scatter plot” comparing a pair of beams (e.g., the ON and one of the OFF beams) and is designed to find bursty emission (with time scales 1 minute). The quantities Q4a to Q4f provide statistical measures of the bursts identified by the Q2 quantity. When examining Q4a-f, the ON and one of the OFF time series are compared to each other; for this, we introduce the difference curve Q4f=Q4f(ON)-Q4f(OFF). We then plot this curve against a reference curve computed from 10000 draws of purely Gaussian noise. The post-processing was performed separately over 3 different frequency ranges (15-26 MHz, 15-39 MHz, and 26-29 MHz).
@@ 5.1 Degenerate causes of the non-detection
To determine the true cause of the potential radio variability of the Boo system, it is crucial to conduct an extensive follow-up campaign. To rule out instrumental systematics, multi-site observations are recommended. It is necessary to monitor the planet’s behavior throughout its orbit and the host star’s magnetic cycle to distinguish between conflicting factors. Additionally, measurements of the star’s magnetic field must be taken during the radio follow-up campaign. Currently, there is an ongoing collaborative campaign utilizing NenuFAR to study Boo radio emission and several telescopes (TBL/Neo-Narval and CHFT/ESPaDOnS) to monitor the magnetic field of the host star. The results of this study will be presented in future work.
@@ 6 Conclusions
In this study, we performed follow-up low-frequency beamformed radio observations of the Boötis ( Boo) exoplanetary system with LOFAR. Previous LOFAR observations showed tentative hints of a detection possibly originating from the exoplanetary system (Turner et al. 2021). In the new observations, we do not detect any burst or slow emission from Boo (Figures 3–4). The current cause of the non-detection is degenerate (see Section 5.1) but could be caused by variability in the planetary radio emission. More radio observations (preferably multi-site) are needed that cover the full orbit of the planet multiple times and at different epochs of the stellar magnetic cycle. Near-simultaneous stellar magnetic maps and stellar monitoring are highly encouraged and needed to disentangle the various competing effects. The newly commissioned NenuFAR and LWA-OVLO are scheduled to observe the Boo exoplanetary system and will help investigate the cause of the possible variable emission.
@@ Acknowledgements
P. Zarka acknowledges funding from the ERC N 101020459—Exoradio.

### v2

@@ Abstract
Context. Observing the radio emission from exoplanets is among the most promising methods to detect their magnetic fields and a measurement of an exoplanetary magnetic field will help constrain the planet’s interior structure, star-planet interactions, atmospheric escape and dynamics, and habitability. Recently, circularly polarized bursty and slow emission from the Boötis ( Boo) exoplanetary system was tentatively detected using LOFAR (LOW-Frequency ARray) beamformed observations. If confirmed, this detection will be a major contribution to exoplanet science. However, follow-up observations are required to confirm this detection.
Aims. Here, we present such follow-up observations of the Boo system using LOFAR. These observations cover 70 of the orbital period of Boo b including the orbital phases of the previous tentative detections.
Methods. We used the BOREALIS pipeline to mitigate radio frequency interference and to search for bursty and slowly varying radio signals. BOREALIS was previously used to find the tentative radio signals from Boo.
Results. Our new observations do not show any signs of bursty or slow emission from the Boötis exoplanetary system.
Conclusions. The cause for our non-detection is currently degenerate. It is possible that the tentative radio signals were an unknown instrumental systematic or that we are observing variability in the planetary radio emission due to changes in its host star. More radio data (preferably multi-site) and ancillary observations (e.g. magnetic maps) are required to further investigate the potential radio emission from the Boötis exoplanetary system.
@@
A confirmed direct detection of exoplanetary radio emission has been elusive for decades. Observing the radio emission from exoplanets is among the most promising methods to detect their magnetic fields (Grießmeier 2015) and a measurement of an exoplanetary magnetic field will help constrain the planet’s interior structure, star-planet interactions, and atmospheric escape (e.g. Zarka et al. 2015; Grießmeier 2015; Lazio et al. 2016; Grießmeier 2017; Lazio 2018; Griessmeier 2018; Zarka 2018; Lazio et al. 2019). A magnetic field measurement can help also help break degeneracies in the mass-radius diagram (Lazio et al., 2019). Additionally, atmospheric dynamics may be influenced by the presence of a planetary magnetic field (e.g. Perna et al. 2010a; Rauscher & Menou 2013; Rogers & Komacek 2014; Hindle et al. 2021) and Ohmic dissipation might be one of the factors causing the anomalously large radii of hot Jupiters (e.g. Perna et al. 2010a, b; Knierim et al. 2022). Finally, a magnetic field might be one of the many properties needed on Earth-like exoplanets to sustain their habitability (e.g. Grießmeier et al. 2005, 2015, 2016; Lammer et al. 2009; Kasting 2010; Owen & Adams 2014; Lazio et al. 2010a, 2016; Meadows & Barnes 2018; McIntyre et al. 2019; Gronoff et al. 2020; Green et al. 2021b).
@@
Observations of auroral radio emissions produced via the Cyclotron Maser Instability (CMI) mechanism (Wu & Lee 1979; Zarka 1998; Treumann 2006) have been used to directly measure the magnetic field strengths of some of the magnetized planets in our Solar System (Burke & Franklin 1955; Zarka 1998). CMI radio signals are emitted at the local electron cyclotron frequency in the source region. Hence, the maximum gyrofrequency is directly proportional to the maximum magnetic field near the planetary surface. CMI radio emission is circularly polarized, beamed, and time-variable (e.g. Zarka 1998; Zarka et al. 2004).
@@
Many theoretical and observational studies have focused on studying exoplanetary radio emissions over the past few decades. The reviews by Zarka et al. (2015) and Grießmeier (2015, 2017) cover the subject in great detail. A large body of theoretical work111e.g. de Pater 2000; Farrell et al. 2004; Lazio et al. 2004; Stevens 2005; Grießmeier et al. 2005, 2007a, 2007b; Jardine & Collier Cameron 2008; Vidotto et al. 2010; Hess & Zarka 2011; Nichols 2011, 2012; Vidotto et al. 2012; Saur et al. 2013; See et al. 2015; Vidotto et al. 2015; Nichols & Milan 2016; Fujii et al. 2016; Vidotto & Donati 2017; Weber et al. 2017a, b, 2018; Lynch et al. 2018; Zarka 2018; Wang & Loeb 2019; Kavanagh et al. 2019; Vidotto et al. 2019; Shiohira et al. 2020; de Gasperin et al. 2020; Green et al. 2021b; Sciola et al. 2021; Narang et al. 2021; Hori 2021; Cendes et al. 2022 has been published since the foundational theoretical studies of Zarka et al. (1997), Farrell et al. (1999), Zarka et al. (2001), and Zarka (2007). Some of these studies aim to develop theoretical models that can predict the frequency and intensity of radio emissions observed from Earth. A variety of scaling laws have been developed to estimate the magnetic field strengths of exoplanets (e.g. Farrell et al. 1999; Sánchez-Lavega 2004; Grießmeier et al. 2007b; Christensen et al. 2009; Christensen 2010; Reiners & Christensen 2010). Likewise, the Radio-Magnetic Bode’s law is a common approach for estimating the radio emission intensity of these planets (e.g. Farrell et al. 1999; Lazio et al. 2004; Zarka 2007; Grießmeier et al. 2007b; Grießmeier 2017; Zarka et al. 2018; Zarka 2018).
@@
Over the past few decades, there have been many exoplanet radio observing campaigns, but none of these observations have resulted in an unambiguous detection222e.g. Yantis et al. 1977; Winglee et al. 1986; Zarka et al. 1997; Bastian et al. 2000; Farrell et al. 2003; Ryabov et al. 2004; Shiratori et al. 2006; George & Stevens 2007; Lazio & Farrell 2007; Smith et al. 2009; Lecavelier Des Etangs et al. 2009, 2011; Lazio et al. 2010a, b; Stroe et al. 2012; Hallinan et al. 2013; Murphy et al. 2015; Lynch et al. 2017; Turner et al. 2017a; O’Gorman et al. 2018; Lenc et al. 2018; Lynch et al. 2018; Route 2019; Narang et al. 2021; de Gasperin et al. 2020; Green & Madhusudhan 2021a; Turner et al. 2021; Narang et al. 2021; Cendes et al. 2022; Narang 2022; Narang et al. 2023a, b; Route & Wolszczan 2023; Turner et al. 2023; Shiohira et al. 2024; Narang et al. 2024. There have been a few tentative detections (Lecavelier des Etangs et al., 2013; Sirothia et al., 2014; Vasylieva, 2015; Bastian et al., 2018), but none of them have been confirmed by follow-up observations. The many degenerate reasons for the non-detections are discussed in great detail in the review articles Zarka et al. (2015) and Grießmeier (2015, 2017). Some reasons for the non-detections could be that the observations were not sensitive enough (e.g. Bastian et al. 2000), there was no emission at the observed frequencies due to a low magnetic field strength (e.g. Hallinan et al. 2013), Earth was outside the CMI beaming pattern (Hess & Zarka 2011; Ashtari et al. 2022), and/or the CMI mechanism was not operating (e.g. plasma frequency of the planet’s ionosphere was greater than the cyclotron frequency; Grießmeier et al. 2007b; Weber et al. 2017a, b; Lamy et al. 2018; Weber et al. 2018). By contrast, advancement has been made in detecting free-floating planets near the brown dwarf boundary (Kao et al. 2016, 2018) and stellar emission potentially created by star-planet interactions (e.g. Vedantham et al. 2020; Callingham et al. 2021; Pérez-Torres et al. 2021; Pineda & Villadsen 2023; Trigilio et al. 2023; Blanco-Pozo et al. 2023).
@@ 2 Observations
All of the observations presented here were taken with the Low Band Antenna (LBA) of LOFAR (van Haarlem et al. 2013) using beamformed mode (Stappers et al. 2011). We obtained 8 exoplanet observations (each 8 hours long) for a total of 64 hours, plus 8 observations of the pulsar B0809+74 totaling 82 minutes. Similar to our previous studies (Turner et al. 2017a, T19, T21, Turner et al. 2023), the pulsar observations are used as a calibrator to study systematics in the data and to ensure the reliability of the processing. An example OFF-beam dynamic spectrum is shown in Figure 2. The dates and times of the observations can be found in Table 1. The setup of the observations can be found in Table 4; it is similar to that in T21. Most importantly, we ensure we cover the 15–30 MHz range where T21 detected their tentative signals. We only present the Stokes-V555Similar to T21, we only analyze as defined in equations 9–11 of Section 4.2.1 in T19. in this paper since the tentative detections were seen in Stokes-V (T21).
@@ 3 Data pipeline
We performed the post-processing of the LOFAR data exactly as in T21. A very detailed description of the burst and slow emission post-processing observables can be found in T19 and T21. A brief summary is given below. The post-processing is performed on the absolute value of the corrected Stokes-V data (as defined in T19). For the slowly varying emission, we calculate the Q1a time series (e.g. Figure 3a-b) and Q1b integrated spectrum (e.g. Figure 3c-d) for each beam (ON, OFF1, OFF2, and OFF3). For burst emission, we only use the Q2 and Q4a-f observable quantities. These quantities are designed to find bursty emission (with time scales 1 minute). The Q2 observable is a time series that is created by high-pass filtering the processed dynamic spectrum and integrating over a large frequency range (e.g. 10 MHz) and time resolution (1 second). This integration over a large frequency range is needed to search for faint emission (see e.g. T21). Q2 is calculated separately for each beam (ON, OFF1, OFF2, and OFF3). Q2 can be represented by a “scatter plot” comparing a pair of beams (e.g. the ON and one of the OFF beams, Figure 4a-b). Next, we calculate the Q4a to Q4f statistical measures to provide a statistical view of the entire observation as individual bursts in the Q2 scatter plots are hard to identify. When examining Q4a-f, the ON and one of the OFF time series are compared to each other; for this, we introduce the difference curve Q4f=Q4f(ON)-Q4f(OFF). We then plot this curve against a reference curve computed from 10000 draws of purely Gaussian noise. The post-processing was performed separately over 3 different frequency ranges (15-26 MHz, 15-39 MHz, and 26-29 MHz).
@@ 5.1 Degenerate causes of the non-detection
To determine the true cause of the potential radio variability of the Boo system, it is crucial to conduct an extensive follow-up campaign. To rule out instrumental systematics, multi-site observations are recommended. We also advocate for complimentary imaging observations to follow-up on the Boo tentative detection (T21). A detailed comparison of the advantages and drawbacks between beamformed and imaging observations can be found in Section 6.3 of T21. It is necessary to monitor the planet’s behavior throughout its orbit and the host star’s magnetic cycle to distinguish between conflicting factors. Additionally, measurements of the star’s magnetic field must be taken during the radio follow-up campaign. Currently, there is an ongoing collaborative campaign utilizing NenuFAR to study Boo radio emission and several telescopes (TBL/Neo-Narval and CHFT/ESPaDOnS) to monitor the magnetic field of the host star. The results of this study will be presented in future work.
@@ 6 Conclusions
In this study, we performed follow-up low-frequency beamformed radio observations of the Boötis ( Boo) exoplanetary system with LOFAR. Previous LOFAR observations showed tentative hints of a detection possibly originating from the exoplanetary system (Turner et al. 2021). In the new observations, we do not detect any burst or slow emission from Boo (Figures 3–4). The current cause of the non-detection is degenerate (see Section 5.1) but could be caused by variability in the planetary radio emission. More radio observations (preferably multi-site) are needed that cover the full orbit of the planet multiple times and at different epochs of the stellar magnetic cycle. Near-simultaneous stellar magnetic maps and stellar monitoring are highly encouraged and needed to disentangle the various competing effects. The newly commissioned NenuFAR and LWA-OVRO are scheduled to observe the Boo exoplanetary system and will help investigate the cause of the possible variable emission.
@@ Acknowledgements
P. Zarka acknowledges funding from the ERC N∘ 101020459—Exoradio.
@@
We thank the anonymous referee for their useful and thoughtful comments.